# Computation method for perfectoid purity and perfectoid BCM-regularity

### v4

@@ 6. Case of graded rings
Let $S=\bigoplus_{i\in\mathbb{Q}_{\geq 0}}S_{i}$ be a Noetherian graded ring such that $S_{0}$ is a divisorial valuation ring with $p$ in the maximal ideal of $S_{0}$, $S$ is $p$-torsion free, and $S_{i}=0$ if $i\notin\mathbb{Z}$. We set $\mathfrak{m}:=(p,S_{>0})$, $R:=S_{\mathfrak{m}}$, and $d:=\dim{\overline{R}}$. Then $R$ satisfies the conditions in Notation 5.1. Furthermore, we define $S^{(m)}:=\bigoplus_{i\geq 0}S_{im}$ as graded ring for every $m\in\mathbb{Q}_{>0}$. For a graded $S$-module $M$ and $m\in\mathbb{Q}$, we define a graded $S$-module $M(m)$ by $M(m)_{i}=M_{i+m}$ for all $i\in\mathbb{Q}$. Since $H^{d}_{\mathfrak{m}}(\overline{R})\simeq H^{d}_{\mathfrak{m}}(\overline{S})$, we endow $H^{d}_{\mathfrak{m}}(\overline{R})$ with a graded structure. We set $a(\overline{S}):=\max\{m\in\mathbb{Z}\mid H^{d}_{\mathfrak{m}}(\overline{S})_{m}\neq 0\}$. Furthermore, we define the $\mathbb{Q}$-graded structure on $W_{n}(S)$ by
@@ 6.1. Case of $a(\overline{S})<0$
We use Notation 6.1. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ be a homogeneous element of degree $\nu$. Then we can run a program in Definition 5.4 such that if we obtain $\eta_{1},\ldots,\eta_{i}$, then there exist $0=e_{0}<e_{1}<\cdots<e_{i}$ such that each $\eta_{j}\in H^{d}_{\mathfrak{m}}(\overline{R})$ is homogeneous of degree $p^{e_{j}}\nu$ for $i\geq j\geq 0$.
@@
We note that $\Phi_{S,n}\colon S\to Q_{S,n}$, $V^{n-1}\colon F^{n}_{*}\overline{S}^{(p^{n})}\to Q_{S,n}$ and $F^{e}\colon\overline{S}\to F^{e}_{*}\overline{S}^{(p^{e})}$ are homomorphisms of graded $S$-modules. We assume that we get a homogeneous element $\eta_{j-1}$ of degree $p^{e_{j-1}}\nu$ by a program in Definition 5.4 for an integer $j\geq 1$. If $\eta_{j-1}$ is in the case (i-i), then $\eta_{j}=F^{e}(\eta_{j})$ for some $e\geq 1$, then $\eta_{j}$ is homogeneous of degree $p^{e_{j-1}+e}\nu$, thus $e_{j}:=e_{j-1}+e$ satisfies the desired condition. If $\eta_{j-1}$ is in the case (ii-i), then there exist an integer $n\geq 2$ and a homogeneous element $\eta^{\prime}$ such that $\Phi_{R,n}(\eta_{j-1})=V^{n-1}(F^{n}_{*}\eta^{\prime})$. Then $\eta^{\prime}$ is of degree $p^{e_{j-1}+n}\nu$, thus $e_{j}:=e_{j-1}+n$ and $\eta_{j}:=\eta^{\prime}$ satisfy the desired condition. ∎
@@
(Theorem B) We use Notation 6.1. We assume $\overline{S}$ is normal quasi-Gorenstein and satisfies $a(\overline{S})<0$. We further assume that $\operatorname{Spec}{\overline{S}}\backslash\{\mathfrak{m}\overline{S}\}$ is strongly $F$-regular. If $R$ is quasi-$F$-split, then $R$ is perfectoid BCM-regular. Furthermore, if $p=2$, the converse implication holds.
@@
We take a perfectoid BCM-algebra $B$ over $R$. First, let us show that $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$ is injective. We take a homogeneous element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ contained in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. Since $a(\overline{S})<0$, the degree of $\eta$ is negative. It is enough to show that $\eta$ is not contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$. Suppose $\eta$ is contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$. We run a program as in Proposition 6.2 and get $\eta_{0}:=\eta,\eta_{1},\ldots,\eta_{i}$. Since $R$ is quasi-$F$-split, $\eta_{i}$ is not in the case (ii-ii). By Lemma 5.6, the element $\eta_{i}$ is contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$. In particular, we have $\eta_{i}\in 0^{*}_{H^{d}_{\mathfrak{m}}(\overline{R})}$ by the proof of [bcm-reg]*Proposition 5.3. By Theorem B.1, there exists a positive integer $e^{\prime}$ such that $F^{e^{\prime}}(\eta_{i})=0$, and in particular, $\eta_{i}$ is not in the case (i-ii). Therefore, by a program, we obtain a sequence of homogeneous elements $\{\eta_{i}\mid i\geq 0\}$ such that there exist $e_{o}=0<e_{1}<\ldots$ with $\deg(\eta_{i})=p^{e_{i}}\deg(\eta)$ for every $i\geq 0$. Since $\overline{R}$ is $F$-pure outside of $\mathfrak{m}\overline{R}$, the kernel of $F\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(F_{*}\overline{R})$ is of finite length, and in particular, there exists $j\geq 0$ such that $F^{e}(\eta_{j})\neq 0$ for every $e\geq 1$. However, it contradicts the fact that the program does not stop. In conclusion, we obtain the injectivity of $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$. Since $\overline{S}$ is quasi-Gorenstein, so is $\overline{R}$, and in particular, the homomorphism $\overline{R}\to\overline{B}$ is pure. Since $\overline{B}$ is a BCM-algebra over $\overline{R}$, we obtain that $\overline{R}$ is Gorenstein, thus so is $R$. Since $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$ is injective and $B$ is a BCM-algebra over $R$, we obtain the injectivity of $H^{d+1}_{\mathfrak{m}}(R)\to H^{d+1}_{\mathfrak{m}}(B)$, thus the homomorphism $R\to B$ is pure, as desired. Next, we assume $p=2$ and $R$ is BCM-regular. Since $R$ is Gorenstein, $R$ is quasi-$F$-split by Corollary 5.3. ∎
@@
and in particular, $S_{0}$ is divisorial valuation ring with the maximal ideal $(p)$. Furthermore, by [bcm-reg]*Theorem 6.27, every stalk of $\mathcal{O}_{X}$ is perfectoid BCM-regular, and in particular, the sheaf $\mathcal{O}_{X}(-mK_{X})$ is Cohen-Macaulay. Since we have $\mathcal{O}_{X}(-mK_{X})\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-mK_{\overline{X}})$ on the codimension one point of $\overline{X}$ and $\mathcal{O}_{X}(-mK_{X})$ satisfies the condition $S_{3}$ for each integer $m$, we obtain $\mathcal{O}_{X}(-mK_{X})\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-mK_{\overline{X}})$. Since $\overline{X}$ is quasi-$F$-split, we obtain $H^{1}(\overline{X},\mathcal{O}_{\overline{X}}(-mK_{\overline{X}}))=0$ for every integer $m\geq 0$ by [KTTWYY1]*Theorem 3.15. By [hartshorne]*Theorem 12.11, we obtain
@@
Indeed, the cone of $(\overline{X},\overline{\Delta})$ coincides with $\overline{S}$ by (a) and (b), the $a$-invariant $a(\overline{S})$ is negative by (c). By an argument in the proof of Corollary 6.4, we have that $X$ is divisorially Cohen-Macaulay, that is, for all divisorial sheaves on $X$ are Cohen-Macaulay. We set $\Delta:=\sum a_{i}E_{i}$, where each $E_{i}$ is a prime divisor and $a_{i}$ has a form $a_{i}=(m_{i}-1)/m_{i}$ for some integer $m_{i}\geq 2$. Since $(\overline{X},\overline{\Delta})$ is strongly $F$-regular, we have $\lfloor\overline{\Delta}\rfloor=0$, and in particular, we have $E_{i}|_{\overline{X}}$ is a reduced divisor. Therefore, we obtain that $\mathcal{O}_{X}(-m(K_{X}+\Delta))\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta}))$ on codimension one points of $\overline{X}$. Since $X$ is divisorially Cohen-Macaulay, we obtain the condition (a). Next, let us show the condition (b). We take an integer $m\geq 0$ and set $L:=\lfloor-m(K_{X}+\Delta)\rfloor$, which is $\mathbb{Q}$-Cartier since $\overline{X}$ is $\mathbb{Q}$-factorial. Since the Picard rank of $\overline{X}$ is one and $-K_{X}$ is ample, we obtain that $L-K_{X}$ or $-L$ is ample. By [KTTWYY1]*Theorem 5.1, we obtain the condition (b) if $\dim{\overline{X}}\geq 2$. Therefore, it is enough to show the condition (c). By the Serre duality, we obtain
@@
(Theorem C) There exists a Noetherian graded ring $S=\bigoplus_{i\in\mathbb{Z}_{\geq 0}}S_{i}$ such that $S$ satisfies the following properties.
@@ 6.2. Case of $a(\overline{S})\geq 0$
We use Notation 6.1. If $R$ is quasi-$F$-split, then we have $a(R)\leq 0$.
@@
Suppose $a(R)>0$, there exists a homogeneous non-zero element $\widetilde{\eta}\in H^{d+1}_{\mathfrak{m}}(R)$ of positive degree. Replacing $\widetilde{\eta}$ by $p^{s}\widetilde{\eta}$ for some $s\geq 0$, we may assume $\widetilde{\eta}$ is $p$-torsion free. Then there exists a homogeneous element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ of positive degree such that $\widetilde{\eta}=\eta/p$. We assume that we get $\eta_{0}:=\eta,\eta_{1},\ldots,\eta_{i}$ by a program as in Proposition 6.2, then $\eta_{j}$ is homogeneous of positive degree for every $i\geq j\geq 0$. Since $R$ is quasi-$F$-split, $\eta_{i}$ is not in the case (ii-ii). Since the degree of $H^{d}_{\mathfrak{m}}(\overline{R})$ is upper bounded, there exists a positive integer $e$ such that $F^{e}(\eta_{i})=0$, thus $\eta_{i}$ is not in case (i-ii). In conclusion, we get $\{\eta_{i}\mid i\geq 0\}$ by such a program. However, by Proposition 6.2, the degrees of $\eta_{i}$ are strictly increasing, thus it contradicts $\eta_{i}\neq 0$ for every $i\geq 0$, as desired. ∎
@@
We use Notation 6.1. We assume $R$ is Gorenstein and $a(\overline{S})=0$. Then
  1. (1)
    $\operatorname{ht}(R)=\operatorname{ht}(\overline{R})$ and
  2. (2)
    $R$ is quasi-$(F,F^{\infty})$-split if and only if $\overline{R}$ is $F$-pure.
@@
Let us show assertion (1). By Proposition 4.6, we have $\operatorname{ht}(R)\leq\operatorname{ht}(\overline{R})$. We set $n:=\operatorname{ht}(R)$, then it is enough to show that $\overline{R}$ is $n$-quasi-$F$-split. If $n=1$, the assertion follows from $Q_{R,1}\simeq Q_{\overline{R},1}\simeq F_{*}\overline{R}$, thus we assume $n\geq 2$. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a homogeneous non-zero element of degree zero, then $\eta$ is contained in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ by $a(\overline{S})=0$. We consider the following commutative diagram;
@@
Next, let us show assertion (2). If $\overline{R}$ is $F$-pure, then $R$ is quasi-$(F,F^{\infty})$-split by definition. We assume $R$ is quasi-$(F,F^{\infty})$-split and set $n:=\operatorname{ht}(R)$. Suppose $n\geq 2$. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a homogeneous non-zero element of degree zero, then it is contained in the socle. We consider the following diagram;
[…]
By the argument in the proof of (1), there exists a homogeneous element $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(\overline{R})$ of degree zero such that $\Phi_{R,n}(\eta)=V^{n-1}(\eta^{\prime})$. Since $\overline{R}$ is not $F$-pure, if $e\geq 2$, then $F^{e-1}(\eta^{\prime})=0$, and it contradicts the fact that $R$ is quasi-$(F,F^{\infty})$-split. Therefore, we obtain that $n=1$, as desired. ∎
@@
(Theorem D) We use Notation Notation 6.1. We assume $R$ is complete intersection and $a(\overline{S})=0$. If $\operatorname{ht}(\overline{R})=n<\infty$, then $R$ is perfectoid pure and we have
@@
By Proposition 6.8, we obtain $\operatorname{ht}(R)=n$. By Theorem 5.12, the ring $R$ is perfectoid pure. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a non-zero element of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$, then $\deg(\eta)=0$. We consider a program as in Proposition 6.2 and get $\eta_{1}\in H^{d}_{\mathfrak{m}}(\overline{R})$, then $\deg(\eta_{1})=0$, and in particular, there exists $u\in\overline{R}^{\times}$ such that $\eta_{1}=u\eta$. If $\eta$ is in the case (i), then $\overline{R}$ is $F$-pure, thus the assertion follows from Theorem 5.8 and Proposition 6.8. We assume $\eta$ is in the case (ii). Since $\operatorname{ht}(R)=n$, we have $\Phi_{R,n}(\eta)=V^{n-1}(F^{n}_{*}\eta_{1})$. We take a perfectoid $R_{\infty}$ as in the proof of Theorem 5.8. We take a system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$ in $R_{\infty}$ and $T\in R_{\infty}^{\flat}$ is the corresponding element of the projective system $\{p^{1/p^{e}}\}_{e\geq 0}$. We set $\tau:=(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))(\eta)$ and $\eta^{\prime}_{1}:=(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))(\tau_{1})$. By the proof of Theorem 5.8, we have
(6.1) $p^{(1-i_{n})}(\tau/p)=wF^{-n}(\tau)/(p^{1/p^{n}}),$
where $w$ is an unit of $R_{\infty}$ and $i_{n}:=1/p+\cdots+1/p^{n}$. We set $\alpha:=\operatorname{ppt}(R;{\rm div}(p))\geq 0$. We consider the diagram
[…]
where $\mathfrak{m}^{\flat}$ is the ideal of $R_{\infty}^{\flat}$ containing $T$ and $\mathfrak{m}^{\flat}/(T)\simeq\mathfrak{m}/(p)$, then the corresponding element to $\tau$ is denoted by $\tau^{\flat}\in H^{d+1}_{\mathfrak{m}}(R_{\infty}^{\flat})$. By (6.1), we have $T^{(1-i_{n})}\tau^{\flat}=w^{\prime}F^{-n}(\tau^{\flat})$ for some unit $w^{\prime}\in R_{\infty}^{\flat}$. Since $F^{-n}$ is an isomorphism, we have $T^{\alpha}\tau^{\flat}=0$ if and only if $T^{1-i_{n}+\alpha/p^{n}}\tau^{\flat}=0$ for $\alpha\in\mathbb{Z}[1/p]_{\geq 0}$. We set $\varepsilon:=\operatorname{ppt}(R;{\rm div}(p))$, then for $\alpha\in\mathbb{Z}[1/p]_{\geq 0}$, we have
$\begin{cases}T^{\alpha}\tau^{\flat}=0&\alpha>\operatorname{ppt}(R;{\rm div}(p))\\ T^{\alpha}\tau^{\flat}\neq 0&\alpha<\operatorname{ppt}(R;{\rm div}(p)).\end{cases}$
Suppose $\operatorname{ppt}(R;{\rm div}(p))>1-(p+\cdots+p^{n-1})/(p^{n}-1)=:\varepsilon_{0}$, then there exists $\alpha\in\mathbb{Z}[1/p]$ such that $\operatorname{ppt}(R;{\rm div}(p))>\alpha>\varepsilon_{0}$. We define a sequence $\{\alpha_{m}\}_{m\geq 1}$ by $\alpha_{1}=\alpha$ and
$\alpha_{m+1}=p^{n}(\alpha_{m}-1+i_{n}),$
then we have $1-i_{n}+\alpha_{m+1}/p^{n}=\alpha_{m}$, and in particular, we have $T^{\alpha_{m}}\tau^{\flat}=0$ for every $m\geq 1$. Since we have $(1-1/p^{n})\varepsilon_{0}=1-i_{n}$, we obtain $\alpha_{m+1}-\varepsilon_{0}=p^{n}(\alpha_{m}-\varepsilon_{0})$. Since $\alpha>\varepsilon_{0}$, there exists $m\geq 1$ such that $\alpha_{m}>\operatorname{ppt}(R;{\rm div}(p))$, thus we get a contradiction. Next, we suppose $\operatorname{ppt}(R;{\rm div}(p))<\varepsilon_{0}$, then by the same argument as above, we obtain a contradiction. Therefore, we have $\operatorname{ppt}(R;{\rm div}(p))=\varepsilon_{0}$, as desired. ∎

### v5

@@ 6. Case of graded rings
Let $S=\bigoplus_{i\in\mathbb{Q}_{\geq 0}}S_{i}$ be a Noetherian graded ring such that $S_{0}$ is a divisorial valuation ring with $p$ in the maximal ideal of $S_{0}$, $S$ is $p$-torsion free, and $S_{i}=0$ if $i\notin\mathbb{Z}$. We set $\mathfrak{m}:=(p,S_{>0})$, $R:=S_{\mathfrak{m}}$, and $d:=\dim{\overline{R}}$. Then $R$ satisfies conditions in Notation 5.1. Furthermore, we define $S^{(m)}:=\bigoplus_{i\geq 0}S_{im}$ as graded ring for every $m\in\mathbb{Q}_{>0}$. For a graded $S$-module $M$ and $m\in\mathbb{Q}$, we define a graded $S$-module $M(m)$ by $M(m)_{i}=M_{i+m}$ for all $i\in\mathbb{Q}$. Since $H^{d}_{\mathfrak{m}}(\overline{R})\simeq H^{d}_{\mathfrak{m}}(\overline{S})$, we endow $H^{d}_{\mathfrak{m}}(\overline{R})$ with a graded structure. We set $a(\overline{S}):=\max\{m\in\mathbb{Z}\mid H^{d}_{\mathfrak{m}}(\overline{S})_{m}\neq 0\}$. Furthermore, we define the $\mathbb{Q}$-graded structure on $W_{n}(S)$ by
@@ 6.1. Case of $a(\overline{S})<0$
We use the notation introduced in Notation 6.1 and assume $R$ is quasi-$F$-split. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ be a homogeneous element of degree $\nu$. Then we can construct sequences $\{\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}\}_{i\geq 0}$ and $\{n_{i}\in\mathbb{Z}_{\geq 1}\}_{i\geq 0}$ using a procedure in Definition 5.4 such that $\eta_{i}$ is a homogeneous element of degree $p^{\sum_{j=0}^{i-1}n_{j}}\nu$.
@@
We note that $\Phi_{S,n}\colon S\to Q_{S,n}$, $V^{n-1}\colon F^{n}_{*}\overline{S}^{(p^{n})}\to Q_{S,n}$ and $F\colon\overline{S}\to F^{e}_{*}\overline{S}^{(p)}$ are homomorphisms of graded $S$-modules. We assume that we get a homogeneous element $\eta_{i}$ of degree $p^{e_{i}}\nu$ by a procedure in Definition 5.4 for an integer $i\geq 0$, where $e_{i}=\sum_{j=0}^{i-1}n_{j}$. If $n_{i}=1$, then $F(\eta_{i})=\eta_{i+1}$ is a homogeneous element of degree $p^{e_{i}+1}\nu$, as desired. If $n_{i}\geq 2$, then since $\Phi_{R,n_{i}}(\eta_{i})$ is homogeneous, we can take $\eta_{i+1}$ as a homogeneous element. Furthermore, the degree of $\eta_{i+1}$ is $p^{e_{i}+n_{i}}\nu$, as desired. ∎
@@
(Theorem C) We use the notation introduced in Notation 6.1. We assume $\overline{S}$ is normal quasi-Gorenstein and satisfies $a(\overline{S})<0$. We further assume that $\operatorname{Spec}{\overline{S}}\backslash\{\mathfrak{m}\overline{S}\}$ is strongly $F$-regular. If $R$ is quasi-$F$-split, then $R$ is perfectoid BCM-regular. Furthermore, if $p=2$, the converse implication holds.
@@
We take a perfectoid BCM-algebra $B$ over $R$. First, we prove that $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$ is injective. We take a homogeneous element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ contained in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. Since $a(\overline{S})<0$, the degree of $\eta$ is negative. By Proposition 6.2, we can take sequences $\{\eta_{i}\}_{i\geq 0}$ and $\{n_{i}\}_{i\geq 0}$ using a procedure in Definition 5.4 such that $\eta_{i}$ is homogeneous of degree $\nu_{i}:=p^{\sum_{j=0}^{i-1}}\nu$. We set
$\eta_{i}^{\prime}:=\left(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})\right)(\eta_{i}).$
Since $\overline{R}$ is $F$-pure outside of $\mathfrak{m}\overline{R}$, the kernel of $F\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(F_{*}\overline{R})$ is of finite length, and in particular, there exists $i\geq 0$ such that $F^{e}(\eta_{i})\neq 0$ for every $e\geq 1$. By Theorem B.1, we have
$\eta_{i}\notin 0^{*}_{H^{d}_{\mathfrak{m}}(\overline{R})}.$
Furthermore, since we have
$\mathrm{Ker}\left(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})\right)\subseteq 0^{*}_{H^{d}_{\mathfrak{m}}(\overline{R})}$
by the proof of [bcm-reg]*Proposition 5.3, we obtain $\eta_{i}^{\prime}\neq 0$. Since $H^{d-1}_{\mathfrak{m}}(\overline{B})=0$, we obtain $\eta_{j}^{\prime}\neq 0$ for every $j\leq i$, inductively. In particular, we have $\eta^{\prime}_{0}\neq 0$ and the map $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$ is injective, as desired.
Since $\overline{S}$ is quasi-Gorenstein, so is $\overline{R}$, and in particular, the homomorphism $\overline{R}\to\overline{B}$ is pure. Since $\overline{B}$ is a BCM-algebra over $\overline{R}$, we obtain that $\overline{R}$ is Gorenstein, thus so is $R$. Since $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$ is injective and $B$ is a BCM-algebra over $R$, we obtain the injectivity of $H^{d+1}_{\mathfrak{m}}(R)\to H^{d+1}_{\mathfrak{m}}(B)$, thus the homomorphism $R\to B$ is pure, as desired. Next, we assume $p=2$ and $R$ is BCM-regular. Since $R$ is Gorenstein, $R$ is quasi-$F$-split by Corollary 5.3. ∎
@@
and in particular, $S_{0}$ is divisorial valuation ring with the maximal ideal $(p)$. Furthermore, by [bcm-reg]*Theorem 6.27, every stalk of $\mathcal{O}_{X}$ is perfectoid BCM-regular, and in particular, the sheaf $\mathcal{O}_{X}(-mK_{X})$ is Cohen-Macaulay. Since we have $\mathcal{O}_{X}(-mK_{X})\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-mK_{\overline{X}})$ on the codimension one point of $\overline{X}$ and $\mathcal{O}_{X}(-mK_{X})$ satisfies condition $S_{3}$ for each integer $m$, we obtain $\mathcal{O}_{X}(-mK_{X})\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-mK_{\overline{X}})$. Since $\overline{X}$ is quasi-$F$-split, we obtain $H^{1}(\overline{X},\mathcal{O}_{\overline{X}}(-mK_{\overline{X}}))=0$ for every integer $m\geq 0$ by [KTTWYY1]*Theorem 3.15. By [hartshorne]*Theorem 12.11, we obtain
@@
Indeed, the cone of $(\overline{X},\overline{\Delta})$ coincides with $\overline{S}$ by (a) and (b), the $a$-invariant $a(\overline{S})$ is negative by (c). By an argument in the proof of Corollary 6.4, we have that $X$ is divisorially Cohen-Macaulay, that is, for all divisorial sheaves on $X$ are Cohen-Macaulay. We set $\Delta:=\sum a_{i}E_{i}$, where each $E_{i}$ is a prime divisor and $a_{i}$ has a form $a_{i}=(m_{i}-1)/m_{i}$ for some integer $m_{i}\geq 2$. Since $(\overline{X},\overline{\Delta})$ is strongly $F$-regular, we have $\lfloor\overline{\Delta}\rfloor=0$, and in particular, we have $E_{i}|_{\overline{X}}$ is a reduced divisor. Therefore, we obtain that $\mathcal{O}_{X}(-m(K_{X}+\Delta))\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta}))$ on codimension one points of $\overline{X}$. Since $X$ is divisorially Cohen-Macaulay, we obtain condition (a). Next, we prove the holding condition (b). We take an integer $m\geq 0$ and set $L:=\lfloor-m(K_{X}+\Delta)\rfloor$, which is $\mathbb{Q}$-Cartier since $\overline{X}$ is $\mathbb{Q}$-factorial. Since the Picard rank of $\overline{X}$ is one and $-K_{X}$ is ample, we obtain that $L-K_{X}$ or $-L$ is ample. By [KTTWYY1]*Theorem 5.1, we obtain condition (b) if $\dim{\overline{X}}\geq 2$. Therefore, it is enough to show condition (c). By the Serre duality, we obtain
@@
(Theorem D) There exists a Noetherian graded ring $S=\bigoplus_{i\in\mathbb{Z}_{\geq 0}}S_{i}$ such that $S$ satisfies the following properties.
@@ 6.2. Case of $a(\overline{S})\geq 0$
We use the notation introduced in Notation 6.1. If $R$ is quasi-$F$-split, then we have $a(R)\leq 0$.
@@
Suppose $a(R)>0$ and $R$ is quasi-$F$-split, then there exists a homogeneous non-zero element $\widetilde{\eta}\in H^{d+1}_{\mathfrak{m}}(R)$ of positive degree. Replacing $\widetilde{\eta}$ by $p^{s}\widetilde{\eta}$ for some $s\geq 0$, we may assume $\widetilde{\eta}$ is $p$-torsion free. Then there exists a homogeneous element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ of positive degree such that $\widetilde{\eta}=\eta/p$. We take a sequence $\{\eta_{i}\}_{i\geq 0}$ as in Proposition 6.2, then the degrees of $\eta_{i}$ are strictly increasing, thus it contradicts $\eta_{i}\neq 0$ for every $i\geq 0$, as desired. ∎
@@
We use the notation introduced in Notation 6.1. We assume $R$ is Gorenstein and $a(\overline{S})=0$. Then $\operatorname{ht}(R)=\operatorname{ht}(\overline{R})$
@@
By Proposition 4.6, we have $\operatorname{ht}(R)\leq\operatorname{ht}(\overline{R})$. We set $n:=\operatorname{ht}(R)$, then it is enough to show that $\overline{R}$ is $n$-quasi-$F$-split. If $n=1$, the assertion follows from $Q_{R,1}\simeq Q_{\overline{R},1}\simeq F_{*}\overline{R}$, thus we assume $n\geq 2$. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a homogeneous non-zero element of degree zero, then $\eta$ is contained in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ by $a(\overline{S})=0$. We consider the following commutative diagram;
@@
(Theorem E) We use Notation Notation 6.1. We assume $R$ is complete intersection and $a(\overline{S})=0$. If $\operatorname{ht}(\overline{R})=n<\infty$, then $R$ is perfectoid pure and we have
@@
By Proposition 6.8, we obtain $\operatorname{ht}(R)=n$. By Theorem 5.7, the ring $R$ is perfectoid pure. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a non-zero element of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$, then $\deg(\eta)=0$. Using a procedure as in Proposition 6.2, we construct a sequence $\{\eta_{i}\}_{i\geq 0}$ with $\eta_{0}=\eta$ and $\{n_{i}\}_{i\geq 0}$, then the degree of $\eta_{i}$ is zero for every $i\geq 0$, and in particular, each $\eta_{i}$ is an element of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. Since $\operatorname{ht}(R)=n$, we have $n_{i}=n$ for every $i\geq 0$. By Theorem 5.9, we have
$\operatorname{ppt}(R;{\rm div}(p))=1-\frac{p+\cdots+p^{n-1}}{p^{n}-1},$
as desired. ∎